### v2

# Scalable Dynamic Embedding Size Search for Streaming Recommendation

Abstract.
Recommender systems typically represent users and items by learning their embeddings, which are usually set to uniform dimensions and dominate the model parameters. However, real-world recommender systems often operate in streaming recommendation scenarios, where the number of users and items continues to grow, leading to substantial storage resource consumption for these embeddings. Although a few methods attempt to mitigate this by employing embedding size search strategies to assign different embedding dimensions in streaming recommendations, they assume that the embedding size grows with the frequency of users/items, which eventually still exceeds the predefined memory budget over time. To address this issue, this paper proposes to learn Scalable Lightweight Embeddings for streaming recommendation, called SCALL, which can adaptively adjust the embedding sizes of users/items within a given memory budget over time. Specifically, we propose to sample embedding sizes from a probabilistic distribution, with the guarantee to meet any predefined memory budget. By fixing the memory budget, the proposed embedding size sampling strategy can increase and decrease the embedding sizes in accordance to the frequency of the corresponding users or items. Furthermore, we develop a reinforcement learning-based search paradigm that models each state with mean pooling to keep the length of the state vectors fixed, invariant to the changing number of users and items. As a result, the proposed method can provide embedding sizes to unseen users and items. Comprehensive empirical evaluations on two public datasets affirm the advantageous effectiveness of our proposed method.
recommender system, reinforcement learning, neural architecture search
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679638isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
1. Introduction
Recommender Systems are algorithms designed to predict and suggest items or content that users might find appealing based on their preferences and behavior (Zhang et al., 2019b; Wang et al., 2021). These systems play a crucial role in various applications, such as online streaming platforms, e-commerce websites, and social media, by enhancing user experience and engagement (Li and Karahanna, 2015; Zhou et al., 2018; Kywe et al., 2012). However, a significant challenge in the implementation of recommender systems lies in the embedding tables, which are used to represent users and items in a low-dimensional embedding vectors of fixed and uniform sizes (Joglekar et al., 2020). As the number of users and items grows in modern applications, the size of the embedding tables becomes a performance and memory bottleneck. For instance, considering Instagram with 2.4 billion monthly active users and 40 billion photos/videos, a conventional embedding table for the users and videos alone would require over 21,708 GB of memory. Managing and processing large embedding tables can lead to increased computational complexity, longer response times, and higher memory requirements, posing a challenge to the efficient operation of recommender systems in resource-constrained environments.
Many researchers have delved into mixed-dimension embedding size strategies aimed at minimizing memory usage while preserving the performance of recommender systems (Qu et al., 2023; Lyu et al., 2022; Qu et al., 2022; Liang et al., 2023; Chen et al., 2021; Qu et al., 2024; Liu et al., 2021; Joglekar et al., 2020). These strategies involve assigning different embedding sizes based on the characteristics of users and items, such as their interaction counts. However, these methods are tailored for static and non-streaming settings and the embedding sizes remain fixed after allotment, which largely limits their practical utility in streaming settings for two reasons. Firstly, in streaming environments, where users and items exhibit high dynamism, the embedding sizes need to be continually adjusted to accommodate the evolving user and item behavior. For example, the embedding size of a newly released movie may need to be increased as its frequency rises over time. Maintaining consistent embedding sizes in such a situation will lead to performance degradation because it fails to adopt to the changing frequency levels. Secondly, these methods cannot to assign embedding sizes to newly joined users or items. As a result, implementing these methods in streaming scenarios necessitates retraining them from scratch at the onset of each time segment, introducing significant computational overheads.
In this regard, several embedding size search mechanisms have emerged to address the challenges of adaptability and efficiency in the evolving landscape of recommendation applications, particularly in streaming scenarios. For instance, Liu et al. (Zhao et al., 2020) proposed AutoEmb, a framework treating the embedding size search as a multi-class classification problem and utilizing a neural network for soft selection of embedding sizes for users and items. The final output embedding size may vary according to the input frequency. Similarly, ESAPN (Liu et al., 2020) and DESS (He et al., 2023) formulate the problem of embedding size adjustment as a binary classification problem. They employ external policies to determine whether to enlarge the embedding size for each user and item as data streams in. However, all of these frameworks lack precise control over their parameter usage, making it challenging to set predefined memory budgets that is crucial for resource-constrained applications. Although BET (Qu et al., 2024) partially addresses this in static recommendation systems using table-level search, which computes the embedding sizes for all users and items, it falls short in streaming settings due to its inability to allocate embedding sizes to newly introduced users and items.
Another problem with ESAPN and DESS is their limited flexibility in adjusting the embedding sizes of previous users and items. Unlike AutoEmb, these methods can only increase embedding sizes for users and items that become more popular, but they cannot reduce them when frequency wanes. Assigning large embedding sizes to users or items that were only popular in the past is unnecessary and inefficient in terms of memory usage. In summary, it is essential to develop a method for embedding size search in streaming recommender systems that have the following advantages: (1) Dynamic Sizes: it can automatically adjust (increase or decrease) embedding sizes to adopt to the changing frequency in each time segment without the need to retrain, (2) Controllable Budget: it can easily scale the embedding table to meet any predefined parameter budget without extra hyperparameter tuning, and (3) Streaming Friendly: it can be easily implemented in streaming settings without retraining in each time segment.
To resolve the three challenges, we propose Scalable Lightweight Embeddings (SCALL) for streaming recommendation systems, an automated embedding size search policy based on the reinforcement learning algorithm Soft Actor-Critic (SAC) (Haarnoja et al., 2018). SCALL achieves (1) and (3) by leveraging the reservoir sampling mechanism of SAC, which enables continual policy learning without the need for periodic retraining. SCALL samples embedding sizes from a probabilistic distribution whose parameters are predicted by the policy network. A controllable memory budget can be effortlessly enforced during the embedding size sampling process, thus achieving (2). Table 1 provides a brief comparison between our proposed method and two other state-of-the-art embedding size search methods, showcasing the superiority of SCALL.
Table 1. Comparison between our method and the most representative state-of-the-art embedding size search methods.
Benefits CIESS BET
ESAPN
DESS
AutoEmb SCALL
Dynamic Sizes
Controllable Budget
Streaming Friendly
Since more frequent users and items usually require higher embedding dimensions, incorporating frequency information into the learning of optimal embedding sizes has been highlighted in several studies (Zhao et al., 2020; Liu et al., 2020; Qu et al., 2023, 2024). However, as the number of users and items increases over time in streaming settings, the size of the state vector becomes variable. To address this issue, we propose the mean pooling strategy, which involves sorting and grouping users and items based on their frequency. Subsequently, we include only the average frequency of each group in each state, ensuring that the length of the state vector remains fixed regardless of the varying number of users and items. Additionally, by sampling embedding sizes from a probabilistic distribution, our proposed method achieves independence from the fluctuating number of users and items, thereby fulfilling (3). The contributions of this paper are outlined below:
  • We pinpoint the practical bottlenecks of current embedding size search methods for streaming recommender systems, specifically their inability to adaptively adjust embedding sizes to conform to a predefined parameter budget.
  • We introduce SCALL, an innovative reinforcement learning (RL)-based approach that dynamically allocates parameters to all the users and items with the guarantee to meet any predefined memory budget in streaming settings, regardless of the changing number of users and items.
  • We improve the table-level search paradigm to accommodate the varying number of users and items in streaming recommendation systems and use it to improve search efficiency.
  • We conduct comprehensive empirical evaluations on two public datasets and compare SCALL to various embedding sparsification algorithms, affirming the advantageous efficacy of SCALL.
2. Relative Work
Before delving into our proposed method, we present the pertinent research topics to offer an in-depth research context from three perspectives: deep recommender systems, streaming recommender systems, and embedding size search in recommendation.
2.1. Deep Recommender Systems
Researchers have put forth a variety of deep recommender models aimed at capturing user-item relations. The initial wave of deep recommender systems is rooted in Multi-Layer Perceptrons (MLPs) (Cheng et al., 2016; He et al., 2017). Cheng et al. (Cheng et al., 2016) introduced Wide&Deep, a model that combines a wide linear model with a deep neural network by computing the weighted sum of their output to leverage the benefits of both memorization and generalization. Another example is Neural Collaborative Filtering (He et al., 2017), which integrates an MLP with matrix factorization to grasp the nuanced user-item interactions. Beyond MLP-based methods, researchers have explored graph-based architectures (Wang et al., 2019; He et al., 2020; Yu et al., 2022). For instance, Neural Graph Collaborative Filtering (Wang et al., 2019) expands neural collaborative filtering by incorporating graph convolution networks to model user-item interactions. LightGCN (He et al., 2020) simplifies NGCF by eliminating self-connections, feature transformations, and linear activation operations. UltraGCN (Mao et al., 2021) further improves LightGCN by using a constraint loss to approximate the limit of infinite-layer graph convolutions. In addition to models that learn from user-item interactions, other data modalities such as location (Yin et al., 2015; Wang et al., 2020), sentiment (Wang et al., 2017a), activity level (Zhang et al., 2024a), text (Li et al., 2023; Zhang et al., 2024b), images (Wang et al., 2017b), and videos (Lee et al., 2018) have also been exploited for recommendation.
2.2. Streaming Recommender Systems
To adapt to ever-evolving landscape of recommendation applications, online learning algorithms that update recommendation model as new data streams in have been studied for streaming recommender systems. For instance, Wang et al. (Wang et al., 2018b) provided a Stream-centered Probabilistic Matrix Factorization model. Guo et al. (Guo et al., 2019) proposed a Streaming Session-based Recommendation Machine to tackle the challenges introduced by user behaviors and large session data volume in streaming recommendation. Qiu et al. (Qiu et al., 2020) introduced a novel Global Attributed Graph neural network model, which learns better session and user representations. Wang et al. (Wang et al., 2018a) proposed a streaming recommender model to capture long-term stable and short-term interests. It is further enhanced by an adaptive negative sampling framework based on Generative Adversarial Nets. To mitigate the problem of catastrophic forgetting and preserve a sketch of users’ long-term interests, reservoir sampling techniques have also been proposed (Zhang et al., 2019a; Chang et al., 2017; Chen et al., 2013; Qiu et al., 2020). They typically involve storing historical user-item interactions in a reservoir, where interactions are sampled using various strategies to update the recommender model. Although these methods can learn user and item representations from continuous streaming data, they still face memory efficiency issues which are yet to be addressed.
2.3. Embedding Compression in Recommendation
A variety of research endeavors have investigated methods for learning compressed embeddings, serving as a memory-efficient alternative to conventional embedding tables in recommender systems, as outlined in (Zheng et al., 2022; Yin et al., 2024; Tran et al., 2024). Approaches like feature selection search, as exemplified by (Wang et al., 2022; Luo et al., 2019; Lin et al., 2022), focus on training models to learn importance scores for effective filtering of unimportant feature fields. Embedding dimension search methods, illustrated by (Liu et al., 2021; Qu et al., 2022; Kong et al., 2023; Qu et al., 2023, 2024), enhance traditional recommender systems by introducing mixed dimensions for features, departing from uniform sizes. Compositional embedding methods (Liang et al., 2024, 2023) represents users and items with a smaller number of embedding vectors (aka. meta-embeddings) to shrink the embedding size. Elastic embedding (Chen et al., 2021; Zheng et al., 2024) enable users and items to share embedding blocks. The knowledge distillation method (Xia et al., 2022, 2023) compresses a model by having a smaller student model learn from a larger teacher model. SCALL maps each user and item to their optimal embedding size so it belong to category of embedding dimension search methods.
Most of the previously mentioned methods offer sparsified embeddings in static recommendation settings. In the context of streaming recommendation tasks, (He et al., 2023; Liu et al., 2020; Zhao et al., 2020) dynamically assign embedding sizes to users and items as data streams in. SCALL is also part of this category, yet it stands out as the first method that combines the three aforementioned advantages: dynamic sizes, controllable budget and streaming friendly.
3. Problem Definition
In this section, we define the task of streaming recommendation and dynamic embedding size prediction.
We create a new data segments each time we witness $m$ new users and items as the data streams in. As a result, the entire data stream is partitioned into $T$ consecutive and mutually exclusive segments arranged in chronological order: $D_{1},\cdots,D_{T-1},D_{T}$. Subsequently, each segment $D_{t}\in{D_{1},\cdots,D_{T-1},D_{T}}$ is further bifurcated into two exclusive parts: the training portion $D_{t}^{tr}$ and the test portion $D_{t}^{te}$. Building upon this framework, the streaming recommendation task is framed as follows: given $D_{1}^{tr},\cdots,D_{T-1}^{tr},D_{T}^{tr}$, the objective is to train a model $G(\cdot)$ to predict the user preferences on $D_{1}^{te},\cdots,D_{T-1}^{te},D_{T}^{te}$. In each time segment $t$, we also build a reservoir, denoted as $\mathcal{R}_{t}$, containing historical interactions that are randomly sampled from $D_{1}\cup...\cup D_{t-2}\cup D_{t-1}^{tr}$ to mitigate catastrophic forgetting. The reservoir sampling strategy is the same as in (Guo et al., 2019).
SCALL operates in a streaming setting. It starts by initializing a base recommender $G_{0}(\cdot)$. The algorithm then iterates through each time segment of the data stream. In each time segment $t\in[1,T]$ with $(D_{t}^{tr}\cup D_{t}^{te})$, the embedding size predictor $F(\cdot)$ is first trained on $D^{tr}_{t-1}$. After the training is completed, $F(\cdot)$ should provide the embedding size for each user and item in $D_{1}\cup...\cup D_{t-1}\cup D_{t}^{tr}$. After obtaining the embedding sizes from the embedding size predictor, the base recommender $G_{t-1}(\cdot)$ should adjust its embedding sizes and update itself to $G_{t}(\cdot)$ on the training samples $D^{tr}_{t}$. After this update, we evaluate $G_{t}(\cdot)$ on $D^{te}_{t}$ and report the recommendation quality w.r.t. Recall@20 and NDCG@20. The algorithm alternates between updating the embedding size predictor and the base recommender in a coordinated manner. The goal of embedding size search is to provide embedding sizes to maximize the recommendation quality on $D^{te}_{t}$ for each time segment $t$. In the remainder of this paper, we omit the subscript $t$ and denote the base recommender as $G(\cdot)$ when there is no confusion.
4. Proposed Methodology
Figure 1. An overview of SCALL.
In a nutshell, SCALL comprises two key components that work alternately: a pre-trained base recommender $G(\cdot)$ parameterized by $\Theta$, and a reinforcement learning-based embedding size predictor $F(\cdot)$ parameterized by $\Phi$. Figure 1 illustrates the workflow of SCALL. In each time segment $t\in[1,T]$, $F(\cdot)$ computes actions that determine the portion of parameters allocated to the users and the parameters of the distributions from which we sample embedding sizes. A fresh base recommender is initialized and adjusts its user/item embedding sizes and tunes its parameters using the training samples from the previous segment $D_{t-1}^{tr}$ with early stopping. Then the base recommender is evaluated on a held-out test set $D_{t-1}^{te}$ w.r.t. Recall@20 and NDCG@20. Based on early stopping Recall@20 and NDCG@20 scores, the embedding size predictor $F(\cdot)$ is revised, and thus the embedding size selection is updated for subsequent iterations. The following sections elucidate the design of SCALL.
4.1. Base Recommender with Adjustable Embedding Sizes
We denote the set of all users as $\mathcal{U}$ and the set of all items $\mathcal{V}$. The recommender’s embedding table maps each user or item to embedding vectors $\mathbf{e}_{n}\in\mathbb{R}^{d_{max}}$, where $n\in\mathcal{U}\cup\mathcal{V}$ and $d_{max}$ is the full embedding size. The vertical concatenation of the embedding vectors for all users and items results in the embedding table $\mathbf{E}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{V}|)\times d_{max}}$. Following the common practice for embedding sparsification in (Liu et al., 2021; Lyu et al., 2022; Qu et al., 2023, 2024), we adjust the embedding sizes through a binary mask matrix $\mathbf{M}\in\{0,1\}^{(|\mathcal{U}|+|\mathcal{V}|)\times d_{max}}$, where each row $\mathbf{m}_{n}\in\mathbf{M}$ is a binary mask vector for user/item $n$, where a proposed embedding size $d_{n}$ sets the first $d_{n}$ elements to ones and the subsequent $d_{max}-d_{n}$ elements to zeros:
(1) $\mathbf{m}_{n}[s]=\begin{cases}1&\text{for }1\leq s\leq d_{n}\\ 0&\text{for }d_{n}<s<d_{max}\end{cases},\;\;\;\;n\in\mathcal{U}\cup\mathcal{V}.$
The embedding table is adjusted during look-up via element-wise multiplication with the binary mask matrix $\mathbf{M}$, i.e., $\mathbf{e}_{n}=(\mathbf{E}\odot\mathbf{M})_{n}$. $\mathbf{M}$ is iteratively updated based on actions given by the policy network. During deployment, only the heavily sparsified embedding table $\mathbf{E}_{sparse}=\mathbf{E}\odot\mathbf{M}$ is stored, where sparse matrix storage techniques (Virtanen et al., 2020; Sedaghati et al., 2015) handle zero-valued entries efficiently. The recommender model $G(\cdot)$ computes the pairwise user-item similarity $\hat{y}_{uv}$ using sparse embedding vectors:
(2) $\hat{y}_{uv}=G(\mathbf{e}_{u},\mathbf{e}_{v}),\;\;\;\;u\in\mathcal{U},v\in\mathcal{V}.$
For optimization, the Bayesian Personalized Ranking loss of the base recommender $G(\cdot)$ on training samples $D^{tr}$ (Rendle et al., 2009) is adopted:
(3) $\mathcal{L}_{BPR}(G(\cdot)|D^{tr})=\sum_{(u,v,v^{\prime})\in\mathcal{D}^{tr}}-\ln\sigma(\hat{y}_{uv}-\hat{y}_{uv^{\prime}})+\eta\|\Theta\|_{2}^{2},$
where $(u,v,v^{\prime})$ denotes user $u$’s preference for item $v$ to $v^{\prime}$, $\Theta$ is the parameter of $G(\cdot)$, and $\eta\|\Theta\|_{2}^{2}$ provides regularization. The goal is to each user and item to their optimal embedding size under a strict memory budget with a fixed mean embedding size: or with a fixed number of parameters, expressed in the overall objective:
(4) $\min_{\Theta,\Phi}\mathcal{L}_{BPR}\;\;\;\text{s.t.}\;\frac{\|\mathbf{M}\|_{1,1}}{(|\mathcal{U}|+|\mathcal{V}|)}\leq c$
where $\|\mathbf{M}\|_{1,1}$ represents the number of ones in the binary mask, $c$ is the desired mean embedding size over all users and items, and $B$ is the desired number of parameters. Alternatively, we can also fix the total number of parameters to $B$:
(5) $\min_{\Theta,\Phi}\mathcal{L}_{BPR}\;\;\;\text{s.t.}\;\|\mathbf{M}\|_{1,1}\leq B$
With fixed $c$, the total number of parameters can be calculated as follows:
(6) $B=c(|\mathcal{U}|+|\mathcal{V}|)$
4.2. Mixed Embedding Size Search with Reinforcement Learning
4.2.1. Environment
The base recommender is the environment. It returns the top-$k$ items for each user given their embedding sizes.
Figure 2. The total parameters are allocated to to 3 users and 3 items in this example. The size of each circle corresponds to the number of parameters, with a larger circle indicating a higher parameter count.
4.2.2. Action
We allocate $wB$ parameters to the users and $(1-w)B$ to the items, where $w\in(0,1)$ denotes the portion of parameters assigned to the users. The users are sorted in descending order of their frequency in $D^{tr}_{t}\cup\mathcal{R}_{t}$. If two users or items have equal frequency, preference is given to the one that appears earlier in the sorted list. Let $\boldsymbol{p}^{U}$ and $\boldsymbol{p}^{V}$ represent the percentage of parameters assigned to each user and item. Inspired by the use of Power Law distribution to model income distribution, we adopt two Power Law distributions, parameterized by $\alpha^{U},\alpha^{V}\in(\alpha_{min},\alpha_{max})$ to separately sample values for $\boldsymbol{p}^{U}$ and $\boldsymbol{p}^{V}$. We can adjust the fairness of parameter allocation by adjusting the values of $\alpha^{U}$ and $\alpha^{V}$. The parameters are uniformly shared by each user/item when $\alpha^{U}$ and $\alpha_{V}$ approach to $\alpha_{max}$; when $\alpha^{U}$ and $\alpha_{V}$ approach to $\alpha_{min}$, the parameters tend to be assigned to the most frequent user/items. In other words, parameter assignment to each user and item is modeled by a mixture model, where each component is a Power Law distribution: $\text{powerlaw}(\alpha)=\alpha x^{\alpha-1}$:
$\boldsymbol{q}^{U}\sim\text{powerlaw}(\alpha^{U});\;\;\;\boldsymbol{q}^{V}\sim\text{powerlaw}(\alpha^{V})$
Subsequently, the distributions are normalized to ensure that the sum of elements in each distribution is equal to 1:
$\boldsymbol{p}^{U}=\frac{\boldsymbol{q}^{U}}{\sum_{u\in|\mathcal{U}|}\boldsymbol{q}_{u}^{U}};\;\;\;\boldsymbol{p}^{V}=\frac{\boldsymbol{q}^{V}}{\sum_{v\in|\mathcal{V}|}\boldsymbol{q}_{v}^{V}}$
Notably, we separate the users and items when sampling their embedding sizes to accommodate the distributional difference in the user and item frequencies. Given a fixed total parameter count, the embedding sizes $\boldsymbol{n}^{U}$ for each user and $\boldsymbol{n}^{V}$ for each item can be easily calculated as follows:
$\boldsymbol{n}^{U}=\lfloor wB\boldsymbol{p}^{U}\rfloor;\;\;\;\boldsymbol{n}^{V}=\lfloor(1-w)B\boldsymbol{p}^{V}\rfloor$
Figure 2 illustrates this parameter allocation process. Clearly, once we decide the values for $w$, $\alpha^{U}$ and $\alpha^{V}$, we can easily derive the embedding size for every user and item. Therefore, we formulate each action as:
(7) $a=(w,\alpha^{U},\alpha^{V})$
4.2.3. State
The input state $s$ is vital in the decision-making process of the policy network by capturing relevant information. Prior research, such as (Zhao et al., 2020; Liu et al., 2020), has highlighted the effectiveness of integrating factors like frequency and contextual details like reward. Following the same design, we integrate the previous reward $r$ in the state.
Figure 3. Mean pooling is used to achieve fixed-length vectors incorporating frequency information. The users or items in this example are divided into 5 groups.
However, we cannot directly incorporate the frequency of every user and item in the state since the number of users and items varies in each data segment. To keep the length of the state fixed, after normalizing the user interaction counts to the range of (0, 1) using minmax normalization, we sort the users in descending order of their frequency and divide them into $L$ groups. Then we calculate the mean frequency of each group and include it in the state. This practice is inspired by the mean pooling operation in Convolution Neural Networks. Figure 3 provides an illustration of this process. The same operation is conducted on the item interaction counts. The resulting pooled user and item frequency is denoted as $\boldsymbol{f}^{U}$ and $\boldsymbol{f}^{V}$.
Likewise, the resulting embedding sizes $\boldsymbol{n}$ cannot be included in the state as contextual information due to their varying length. In light of this challenge, we only include the range of $\boldsymbol{n}^{U}$ and $\boldsymbol{n}^{V}$, normalized by $d_{max}-d_{min}$, to measure their dispersion:
$h^{U}=\frac{\text{max}(\boldsymbol{n}^{U})-\text{min}(\boldsymbol{n}^{U})}{d_{max}-d_{min}};\;\;\;h^{V}=\frac{\text{max}(\boldsymbol{n}^{V})-\text{min}(\boldsymbol{n}^{V})}{d_{max}-d_{min}}$
It is worth noting that variance or standard deviation cannot be used here because they are sensitive to the scale of the data. Although $\alpha^{U}$ or $\alpha^{V}$ can measure the statistical dispersion of $\boldsymbol{n}$, they are not suitable for use here because they are invariant of the data scale, which can be exemplified by the case of $\boldsymbol{n}_{1}=[128,64,64]$, $\boldsymbol{n}_{2}=[2,1,1]$ $\boldsymbol{n}_{3}=[1,1,1]$. With the use of $\alpha^{U/V}$, $\boldsymbol{n}_{1}$ and $\boldsymbol{n}_{2}$ will be deemed as equivalent, while in reality $\boldsymbol{n}_{2}$ makes little difference from $\boldsymbol{n}_{3}$ in affecting the performance of the base recommender when they are used as embedding sizes.
Moreover, to capture the characteristics of the data segment, the mean user/item embeddings (denoted as $\boldsymbol{m}^{U},\boldsymbol{m}^{V}$) and the user number to item number ratio $\frac{|\mathcal{U}|}{|\mathcal{U}|+|\mathcal{V}|}$ are also included by the state. The mean embeddings are calculated as follows:
$\boldsymbol{m}^{U}=\frac{\sum_{u\in\mathcal{U}}(\mathbf{E}\odot\mathbf{M})_{u}}{|\mathcal{U}|};\;\;\;\boldsymbol{m}^{V}=\frac{\sum_{v\in\mathcal{V}}(\mathbf{E}\odot\mathbf{M})_{v}}{|\mathcal{V}|}$
Similar to the computation of $\boldsymbol{f}^{U}$ and $\boldsymbol{f}^{V}$, we take the mean of the user and item embedding sizes because their vector length is independent of the number of users and items.
In summary, our method constructs the state $s$ as follows:
(8) $s=(\boldsymbol{f}^{U},\boldsymbol{f}^{V},r,\frac{|\mathcal{U}|}{|\mathcal{U}|+|\mathcal{V}|},h^{U},h^{V},\boldsymbol{m}^{U},\boldsymbol{m}^{V})$
4.2.4. Reward
In each iteration, the base recommender dynamically adjusts its embedding sizes based on the action $a$. Subsequently, a reward $r$ is generated to provide feedback to the current policy and guide the following embedding size allocations. Aligned with our objective equations Eq.(4) and Eq. (6), the reward needs to effectively capture the recommendation quality. Building on the findings in (White et al., 2021), which highlight the effectiveness of accuracy with early stopping as a performance predictor in neural architecture search, we utilize Recall@20 and NDCG@20 with early stopping to formulate the reward. Specifically, a fresh base recommender is initialized for fine-tuning before its recommendation quality is evaluated. We initialize a new recommender and do not inherit any pre-trained embeddings so that the previous embedding sizes do not have any influence on the ongoing embedding size search. The agent will have the tendency to preserve the previous embedding sizes if we directly fine-tune a trained base recommender. Hence, we define the following reward for an action $a$ in time segment $t$:
(9) $r\propto\frac{\text{eval}(G^{\prime}_{t,j}|\mathcal{D}^{te}_{t})}{\text{eval}(G_{t,j}|\mathcal{D}^{te}_{t})},$
where $r$ is further normalized to fall within the range of $(0,10)$, $G_{t,j}$ represents the snapshot of the base recommender that is updated on $D^{tr}_{t}$ at the $j$-th step, $G^{\prime}_{t,j}$ represents the newly initialized base recommender after being fine-tuned on $D^{tr}_{t}$ for $j$ steps, $\text{eval}(\cdot|\mathcal{D}_{t}^{te})$ is a recommendation quality measure on the test set $\mathcal{D}_{t}^{te}$. It is constructed as an ensemble of Recall@20 and NDCG@20, which are both common recommendation metrics (He et al., 2020; Rendle et al., 2009):
(10) $\text{eval}(\cdot)=\sum_{u\in\mathcal{U}}\frac{\text{Recall@}20_{u}+\text{NDCG@}20_{u}}{2|\mathcal{U}|}$
$G_{1,j}$ is obtained by training a freshly initialized recommender on $D_{1}^{tr}$ with fixed and uniform sizes.
Importantly, Recall@20 and NDCG@20 scores are not directly comparable across different time segments due to the inherent variability in data characteristics. Therefore, instead of using $\text{eval}(G^{\prime}_{t,j}|\mathcal{D}^{te}_{t})$ as the reward, we compare it to $\text{eval}(G_{t,j}|\mathcal{D}^{te}_{t})$ to reflect the performance gain/loss. This scaling ensures comparability across each time segment, allowing it to reflect performance fluctuations with respect to the performance before the embedding size adjustment.
4.2.5. Actor and Critic
SCALL adopts the actor-critic paradigm in reinforcement learning. We employ Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the reinforcement learning backbone. SAC consists of an actor network $\pi(\cdot)$ and a critic network $Q(\cdot)$. In each iteration, the actor network $\pi(\cdot)$ maps current state $s$ to an action $a$:
(11) $a=\pi(s)$
To approximate the quality (Q-value) of an action $a$ taken at state $s$, we employ a critic network denoted as $Q(\cdot)$. The critic networks are updated using transition tuples stored in the replay buffer $\mathcal{Z}$. The complete pseudo-code of SCALL is presented in Algorithm 1.
Algorithm 1 SCALL
1:Initialize replay buffer $\mathcal{Z}$;
2:Initialize base recommender $G_{0}(\cdot)$;
3:for $t=1,\cdots,T$ do
4:     if t ¿ 1 then
5:         for $i=0,\cdots,N$ do
6:              /*  Update embedding size predictor  */
7:              Get last time segment of data $(D^{tr}_{t-1}\cup D^{te}_{t-1})$;
8:              Obtain $a_{i}\leftarrow$ Eq.(11) and temporarily update $\mathbf{M}$;
9:              Tune the newly initialized $G^{\prime}_{t-1}(\cdot)$ on $D^{tr}_{t-1}$;
10:              Calculate reward $r_{i}\!\!\leftarrow\!$ Eq.(9)
11:              Get subsequent state $s’\!\!\leftarrow\!$ ’ ← Eq.(8);
12:              Update buffer $\mathcal{Z}\leftarrow\mathcal{Z}\cup(s_{i},a_{i},r_{i},s^{\prime}_{i})$;
13:              Randomly sample a batch of transitions from $\mathcal{Z}$;
14:              Update $Q(\cdot)$ and $\!\pi(\cdot)$ as in SAC (Haarnoja et al., 2018);
15:         end for
16:     end if
17:     /*  Update base recommender  */
18:     Get current time segment of data $(D^{tr}_{t}\cup D^{te}_{t})$;
19:     if t = 1 then
20:         Update $\mathbf{M}$ with fixed and uniform sizes;
21:     else:
22:         Obtain action $a\leftarrow$ Eq.(11) and permanently update $\mathbf{M}$;
23:     end if
24:     Permanently update $G_{t-1}(\cdot)$ to $G_{t}(\cdot)$ on $D^{tr}_{t}$;
25:     Evaluate $G_{t}(\cdot)$ on $D^{te}_{t}$ w.r.t. Recall@20 and NDCG@20;
26:end for
Table 2. Performance of all methods on Yelp and Amazon-Book with mean embedding size set to 64, 32, and 16. The best results are highlighted. Recall@20 and NDCG@20 are abbreviated to R@20 and N@20 in this and all the subsequent tables for simplicity. All our results are statistically significant with $p<0.02$.
Yelp Amazon-Book
Method c = 16 c = 32 c = 64 c = 16 c = 32 c = 64
R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
ES 0.0468 0.0253 0.0524 0.0281 0.0571 0.0303 0.0411 0.0229 0.0478 0.0265 0.0497 0.0279
MR 0.0453 0.0243 0.0488 0.0260 0.0526 0.0281 0.0376 0.0208 0.0416 0.0229 0.0452 0.0249
CIESS 0.0536 0.0295 0.0564 0.0309 0.0585 0.0319 0.0419 0.0235 0.0484 0.0274 0.0520 0.0293
BET 0.0542 0.0297 0.0581 0.0318 0.0596 0.0324 0.0431 0.0243 0.0501 0.0282 0.0529 0.0300
SCALL 0.0560 0.0305 0.0604 0.0329 0.0623 0.0335 0.0461 0.0258 0.0516 0.0290 0.0554 0.0310
c R@20 N@20 c R@20 N@20
ESAPN 84 0.0592 0.0323 78 0.0529 0.0299
DESS 39 0.0586 0.0318 46 0.0503 0.0285
AutoEmb 74 0.0560 0.0307 65 0.0489 0.0277
Table 3. Performance of all methods on Yelp and Amazon-Book with the parameter budget set to 10MB and 5MB. The best results are highlighted. All our results are statistically significant with $p<0.03$.
Yelp Amazon-Book
Method B = 5MB B = 10MB B = 5MB B = 10MB
R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
ES 0.0474 0.0256 0.0520 0.0279 0.0417 0.0231 0.0475 0.0263
MR 0.0386 0.0210 0.0496 0.0270 0.0312 0.0175 0.0356 0.0196
CIESS 0.0502 0.0278 0.0538 0.0296 0.0414 0.0233 0.0483 0.0274
BET 0.0524 0.0289 0.0557 0.0204 0.0439 0.0248 0.0495 0.0280
SCALL 0.0545 0.0300 0.0579 0.0316 0.0462 0.0259 0.0515 0.0291
B R@20 N@20 B R@20 N@20
ESAPN 9.29MB 0.0539 0.0300 8.12MB 0.0442 0.0248
DESS 7.66MB 0.0520 0.0287 6.30MB 0.0450 0.0252
AutoEmb 7.98MB 0.0505 0.0278 9.76MB 0.0458 0.0258
5. Experiments
In this section, we elucidate our experimental analysis on SCALL. Through the experiments, we seek to answer the following research questions:
  • RQ1: how does SCALL perform compared to ES and other embedding size search methods in streaming settings without the need to be retrained?
  • RQ2: Can SCALL enforce any predefined memory budget during embedding size search?
  • RQ3: Can SCALL dynamically adjust embedding sizes for users and items?
  • RQ4: How does the reservoir size $\gamma$ impact the performance?
  • RQ5: Which probabilistic distribution yields the best performance?
5.1. Baseline Methods
To verify the superiority of the proposed method, we compare SCALL against several embedding size search algorithms that can be classified into three groups: fixed embedding size (i.e., ES), static methods (i.e., MR, CIESS and BET) and dynamic methods (i.e., ESAPN, DESS, AutoEmb and SCALL) for streaming recommendation. In static methods, the user and item embedding sizes remain fixed. We retrain them from scratch in every data segment so they are comparable to dynamic methods, which adjusts the embedding sizes in each time segment. The baseline methods are listed below:
  • CIESS (Qu et al., 2023): An TD3-based embedding size search algorithm that chooses the embedding sizes from a continuous domain.
  • BET (Qu et al., 2024): A neural architecture search method that samples numerous table-level candidate embedding sizes and selects the best action using a DeepSets-based fitness predictor.
  • AutoEmb (Zhao et al., 2020): A differentiable framework that can automatically adjust embedding sizes based on user/item popularity in a streaming setting.
  • ESAPN (Liu et al., 2020): A reinforcement learning-based embedding size search algorithm that performs hard selection on embedding sizes based on memory complexity and recommendation performance.
  • DESS (He et al., 2023): A non-stationary LinUCB-based method with a sublinear regret upper bound and solves the problem of embedding size search by minimizing the cumulative regret.
  • Equal Sizes (ES): ES assigns equal embedding sizes for all users and items.
  • Mixed and Random (MR): MR samples random embedding sizes from a uniform distribution.
5.2. Datasets and Evaluation Protocols
The proposed and baseline method undergo evaluation on two real-world datasets: Amazon-Book (Ni et al., 2019) and Yelp (Wang et al., 2019), both of which contain user ratings. We apply 30-core and 10-core processing on the two datasets, respectively. All the remaining interactions are treated as positive and are sorted in chronological order. After processing, the former dataset comprises 5,076,971 interactions involving 66,351 users and 57,270 items, while the latter involves 2,762,098 interactions between 99,011 users and 56,441 items. A new data segment is created whenever there are 15,000 users or items are introduced (i.e., $m=15,000$). We partition each data segment into 80% for training and 20% for testing, ensuring that the training data precedes the test data chronologically. As a result, 10 data segments are created for the Yelp dataset and 8 data segments are created for the Amazon-Book dataset.
We employ Recall@20 and NDCG@20 as evaluation metrics. The final performance of each embedding sparsification method is evaluated using the mean Recall@20 and NDCG@20 scores over every data segment.
We also test SCALL under two scenarios - fixed mean embedding sizes and fixed numbers of parameters. In the first scenario, we set the mean embedding size $c$ to 16, 32, and 64. In second scenario, the number of parameters $B$ is set to 1,250,000 and 2,500,000, corresponding to 5MB and 10MB memory. It is worth pointing out that memory budgets larger than 10MB (e.g., 15MB) are likely to cause the mean embedding size to exceed 256, which is the full embedding size in our setting, and is thus unnecessary. However, as AutoEmb, ESAPN and DESS are not explicitly designed for precise control over embedding sparsity and prioritize performance, this section solely presents their performance linked to the final embedding tables after pruning.
5.3. Implementation Details
In SCALL, we iteratively tune the base recommender to search for the optimal embedding sizes. Once the policy network converges, we train the base recommender till it converges. LightGCN (He et al., 2020) is employed as the base recommender. During the training and tuning phases, we utilize a batch size of 10,000. The base recommender is updated 200 times each time it is tuned. During the training phase, learning rate decay is applied. The initial learning rate is set to 0.03 and decays every 200 training steps, with 0.95 decay rate. The minimum learning rate set to 0.001. The full embedding size $d_{max}$ is 256 and the minimal embedding size $d_{min}$ is 1. Early stopping is integrated into the training phase, with evaluation occurring every 200 steps of updates and an early stopping patience of 3.
When training the policy network, $\alpha_{max}$ and $\alpha_{min}$ are set to 30 and 0.1 for Power Law distributions, $10\times 10^{-5}$ and $10\times 10^{-6}$ for normal distributions, 100 and 30 for Pareto distributions. The user and item interaction counts are each divided into $L=256$ groups when we calculate $\boldsymbol{f}^{U}$ and $\boldsymbol{f}^{V}$. Both the actor and critic networks are MLPs with one hidden layer and the hidden size is set to 512. 8 transition tuples are sampled from the replay buffer in each iteration.
5.4. Overall Performance Comparison
Table 2 and Table 3 present the Recall@20 and NDCG@20 scores obtained from various embedding sparsification methods with fixed mean embedding sizes and fixed numbers of parameters, respectively. Additionally, we conduct paired Student’s $t$-tests between SCALL and the best-performing baseline method in each setting. As shown in both tables, across sparsity levels ($c\in\{16,32,64\}$ and $B\in\{5\text{MB},10\text{MB}\}$), SCALL consistently outperforms the baselines with adjustable sparsity ratios (i.e., ES, MR, BET, and CIESS) in all settings. MR has the worst performance in all settings. Notably, as shown by Table 2, at a higher sparsification level, SCALL with $c=64$ surpasses ESAPN, DESS and AutoEmb with higher memory usage on both datasets, which proves the superiority of SCALL. Table 3 also shows that, although using more parameters ($B>5\text{MB}$), ESAPN, DESS and AutoEmb do not outperform SCALL with $B=5\text{MB}$. The associated $p$-values, all below 0.05, provide statistical significance, indicating that the advantageous efficacy of SCALL results are not random and thus answering RQ1. Tables 2 and 3 further demonstrate that, unlike ESAPN, DESS and AutoEmb, SCALL can enforce any memory budget, thus addressing RQ2. It’s worth noting that, through tuning the scaling coefficient in its reward function, CIESS also achieves memory budget enforcement (Qu et al., 2023), but this approach is often impractical in real-world implementations.
It is important to highlight that while CIESS and BET demonstrate respectable performance across various settings, their lack of compatibility with streaming recommendation settings entails initialization and training from scratch in each data segment. Consequently, the associated computational costs render them impractical for large-scale recommendation applications.
5.5. Model Component Analysis
5.5.1. User-item Separation
As described in section 4.2.2, we separate users from items and sample their embedding sizes from different distributions due to their distributional difference. To verify the performance benefits stemming from this design choice, we conduct an experiment where we aggregate users and items into a unified set of $L$ groups, ordered by descending frequency in $D_{t}^{tr}\cup\mathcal{R}$. We also sample the embedding sizes from one single Power Law distribution parameterized by $\alpha$. The state and action are simplified as follows:
$s$ $=(\boldsymbol{f},r,\frac{|\mathcal{U}|}{|\mathcal{U}|+|\mathcal{V}|},\boldsymbol{m}^{U},\boldsymbol{m}^{V})$
(12) $a$ $=\alpha,$
where $\boldsymbol{f}$ contains the sorted normalized frequency for all users and items. The mean embedding size $c$ is 32. We refer to this simplified version of SCALL as Dummy and compare its performance against the standard version, which maintains user item separation. The results, presented in Table 5, demonstrate the efficacy of separately grouping users and items in enhancing performance.
5.5.2. State Information
As outlined in Section 4.2.3, the state representation incorporates information about the statistical dispersion of resulting embedding sizes, encapsulated in their normalized range denoted as $h^{U}$ and $h^{V}$ in each state. Additionally, to capture the latent information within each data segment and determine the corresponding embedding sizes, we calculate the mean user embedding vectors $\boldsymbol{m}^{U}$ across all users and the mean item embedding vectors $\boldsymbol{m}^{V}$ across all items. To assess the impact of $h^{U}$ and $h^{V}$ on the final recommendation quality, we exclude them from the state representation and report the Recall@20 and NDCG@20 scores. Similarly, we omit $\boldsymbol{m}^{U}$ and $\boldsymbol{m}^{V}$ from the state representation before algorithm execution and report the same metrics. We set the mean embedding size $c$ to 32. The results are displayed in Table 4. By comparing these results with the performance of the standard version of SCALL, we conclude that removing any of these components results in a degradation of performance.
Table 4. Performance of different variations of SCALL after one of the state components is removed.
Yelp Amazon-Book
Variation
R@20 N@20 R@20 N@20
w/o $\boldsymbol{m}^{U/V}$ 0.0594 0.0324 0.0510 0.0286
w/o $h^{U/V}$ 0.0589 0.0323 0.0507 0.0284
Dummy 0.0501 0.0268 0.0387 0.0217
Table 5. Performance metrics of SCALL variations adopting different distributions.
Yelp Amazon-Book
Distribution R@20 N@20 R@20 N@20
Pareto 0.0602 0.0325 0.0512 0.0287
Normal 0.0602 0.0329 0.0504 0.0285
Power Law 0.0604 0.0329 0.0516 0.0290
5.6. Analysis of Hyperparameters
In this section, we examine the impact of fundamental hyperparameters within SCALL w.r.t. their influence on Recall@20 and NDCG@20.
5.6.1. Reservoir Size $\gamma$
Figure 4. Sensitivity analysis of $\gamma$ w.r.t. Recall@20 and NDCG@20 on the Yelp and Amazon-Book dataset.
During time segment $t$, the base recommender undergoes training on the combined dataset $D^{tr}_{t}\cup\mathcal{R}_{t}$, where $\mathcal{R}_{t}$ represents a reservoir containing historical user-item interactions sampled from $D_{1}\cup...\cup D_{t-2}\cup D_{t-1}$. The reservoir’s size is determined as $\gamma|D^{tr}_{t}|$, where $|D^{tr}_{t}|$ represents the number of interactions in $D^{tr}_{t}$. Users and items are sorted based on their frequency in $D^{tr}_{t}\cup\mathcal{R}_{t}$, allowing for the adjustment of the influence that past records exert on the current embedding size search. To answer RQ4, we explore the impact of different $\gamma$ values from $\{0.1,0.2,0.3,0.4,0.5\}$ on Recall@20 and NDCG@20 scores to investigate how the reservoir size affects recommendation performance. The mean embedding size $c$ is set to 32. The results, depicted in Figure 4, reveal that SCALL achieves optimal performance when $\gamma$ is set to $0.1$ for Amazon-Book. The performance of SCALL is generally insensitive to the setting of $\gamma$ on the Yelp dataset, showcasing no significant difference with varying values of $\gamma$.
5.6.2. Choice of Distributions
As outlined in Section 4.2.2, the distribution of parameters among user and item groups is determined by sampling values from two Power Law distributions. To answer RQ5, we replace the Power Law distributions with truncated normal distributions and Pareto distributions, which are commonly used to model social income and wealth distribution. Similar to the Power Law distributions, both types of distributions are parameterized by $\alpha_{U}$ and $\alpha_{V}$, which is bounded by $(\alpha_{min},\alpha_{max})$. For the normal distributions, they are bounded within $(0,1)$. The normal distribution for the users has mean of $\frac{1}{|\mathcal{U}|}$ and the one for the items has mean of $\frac{1}{|\mathcal{V}|}$. Their scales are set to $\alpha^{U}$ and $\alpha^{V}$. The Pareto distributions are parameterized by $\alpha^{U}$ and $\alpha^{V}$. The fairness of parameter allocation can be adjusted by adjusting the values of $\alpha^{U}$ and $\alpha^{V}$. The parameters tend be to distributed evenly or to the most frequent users and items when $\alpha^{U}$ and $\alpha^{V}$ approach one of the two extremes. The mean embedding size $c$ is set to 32. The performance comparison presented in Table 5 indicates that SCALL is generally insensitive to the choice of distributions, with Power Law and Pareto distributions achieving slightly better performance.
5.7. Case Study of User Embedding Sizes
Figure 5. Case study on the interaction counts (left) and embedding sizes (right) of four users from the Yelp dataset.
To gain insight into how the embedding size varies for specific users or items across different time segments, we focus on four users: user 1546, user 2250, user 3211 and user 10120. We execute the algorithm on the Yelp dataset with the mean embedding size set to 32. Then we generate plots illustrating their interaction counts, which represents their frequency, and corresponding embedding sizes. Each of these users represents a distinct user profile. Users are represented in the plots only after their initial appearance in the dataset. User 1546 consistently maintains high frequency across all segments. In contrast, user 3211 remains unpopular throughout all segments. User 10120 initially starts with low frequency in the first four segments but experiences a steady increase in frequency over the subsequent six segments. Conversely, user 2250 experiences a rise in frequency from the second to the first time segment, followed decline in the subsequent three segments and a slight rise over the seventh and the eighth segments.
As depicted in Figure 5, all four users initially start with large embedding sizes despite their significant differences in frequency. This phenomenon arises because the number of users is relatively small in the initial segments, allowing each user to receive a relatively generous allocation of embedding sizes. However, user 2250 and user 3211 are still given the smallest embedding size, reflecting its lower frequency compared to the others.
Over time, the embedding size of user 1546 remains consistently large due to its sustained frequency. In contrast, the embedding sizes of users 3211 and 10120 begin to lag behind user 1546 in the first three segments, reflecting their lower frequency. However, while user 10120 initially follows a similar trend as user 3211, its embedding size quickly catches up after the third segment due to an increase in frequency. The embedding size of user 2250 increases in the second segment and decreases in the following three segments, which is consistent with the trend of its frequency. This frequency later experiences a slight rise in the seventh and eighth segments, leading to an increase in its embedding size. Comparing the embedding sizes of user 10120 and user 1546 after the fourth segment, we find they share the same embedding size although user 1546 is much more frequent. This reveals that SCALL effectively prevents any individual user from receiving an excessive number of parameters by capping the embedding size of highly frequent users. This conclusion is reinforced by examining the frequency and embedding size patterns of users 1546 and 2250 in the seventh and eighth segments.
Furthermore, despite user 3211 having equal frequency to user 2250 in the last two segments, its embedding size is slightly larger. This discrepancy arises because SCALL prioritizes newer users or items when they have the same frequency. Consequently, user 3211, being a newer addition to the dataset compared to user 2250, receives a larger embedding size allotment.
In conclusion, SCALL is able to dynamically adjust the embedding sizes of users and items according to their changing frequency across different data segments, which answers RQ3.
6. Conclusion and Future Research
Recommender Systems face a significant challenge in the form of the memory bottleneck resulting from uniform embedding sizes for all users and items. Existing dynamic embedding size strategies aim to minimize memory footprint but do not have any fixed constraint on the parameter budget. To address these issues, we proposed SCALL, an automated policy that dynamically adjusts embedding sizes for each user and item and ensures adherence to a predefined budget. The experimental results on the Yelp and Amazon-Book datasets verified the advantageous efficacy of SCALL. However, performing embedding size assignment based on the user and item frequency is biased against the users and items with low frequency. Moving forward, we will concentrate on fair embedding size allocation that overcomes such frequency bias.
Acknowledgment
This work is supported by Australian Research Council under the streams of Future Fellowship (Grant No. FT210100624), Linkage Project (Grant No. LP230200892), Discovery Early Career Researcher Award (Grants No. DE230101033), Discovery Project (Grants No. DP240101108 and No. DP240101814).